An object is being swirled in a circle at the end of a string. At time t = 0, the angular speed of the object is 0.6 rad/s. If the torque on the system produces an angular acceleration of 0.13 rad/s2, what is the angular speed after 25 rad of angular motion?
Answer #1

Given,

w0 = 0.6 rad/s ; alpha = 0.13 rad/s^2 ; theta = 25 rad

We know from eqn of motion in circular motion

w^2 = w^2 + 2 alpha theta

w^2 = 0.6^2 + 2 x 0.13 x 25 = 6.86

w = sqrt(6.86) = 2.62 rad/s

Hence, w = 2.6 rad/s